# Non-Equiprobable Signaling for Wireless Channels Subject to Mobility and Delay Spread

### v1

@@ Abstract
This letter describes how to improve performance of OFDM systems by combining non-equiprobable signaling with low density parity check (LDPC) coding. We partition a standard QAM constellation into annular subconstellations of equal size, and we implement non-equiprobable signaling through a shaping code which selects subconstellations with large average energy less frequently than subconstellations with small average energy. In equiprobable signaling, the LDPC code selects a signal point from the inner subconstellation. In non-equiprobable signaling this inner signal point has a representative in each subconstellation and the shaping code selects the representative for transmission. It is possible to use standard QAM constellations to achieve any desired fractional bit rate with this method of shaping the energy distribution of the transmitted signal. We describe how to combine coding and shaping by integrating shaping into the calculation of log-likelihood ratios (LLRs) necessary for decoding LDPC codes. We present simulation results for non-equiprobable transmission at $1.5$ bits/symbol on a representative Veh-A channel showing gains of $4$ dB at a bit error rate (BER) of $10^{-3}$. As the transmission rate increases, the gains from non-equiprobable signaling diminish, but we show through simulation that they are still significant for $16$-QAM.
@@ I Introduction
In his 1948 paper [shannon1948a], Shannon recognized that signals input to a Gaussian channel should themselves be selected with a Gaussian distribution. Wireless channels subject to mobility and delay spread are not Gaussian channels, but we will show that it is still advantageous to select signals with large energy less frequently than signals with small energy. We will transmit some information through our choice of signal energy, and since signaling is no longer equiprobable, it will be necessary to expand the signal constellation.
Section II describes the OFDM system model. Section III introduces our method of shaping on rings through a small example, where the baseline system is LDPC-coded OFDM with a $4$-QAM signal constellation. We expand to the $8$-QAM constellation shown in Fig. 2, and we use a shaping code to select points from the inner and outer subconstellations with different probabilities. Section IV compares BER performance of non-equiprobable signaling (shaping) with equiprobable signaling (4-QAM) on a representative Veh-A channel [veh_a]. For the same communication rate ($1.5$ bits per $2$-dim. symbol) we present simulations showing a $4$ dB gain of rate $1/2$ LDPC coding + rate $1/2$ shaping over rate $3/4$ LDPC coding. The cost is a modest increase in Peak to Average Power Ratio (PAPR). As the transmission rate increases, the possible shaping gains diminish, but we are still able to present an example in Section IV-C showing significant gains for $16$-QAM. We would also argue that $4$-QAM is the most important modulation and coding scheme.
It is also possible to change the distribution of transmitted signal energy by warping a standard QAM constellation so that points closer to the origin are closer together and points closer to the periphery are further apart [hoydis2020joint]. Shaping on rings makes it possible to achieve bit error rate (BER) improvements with standard QAM signal constellations.
The method of shaping on rings is quite general. We partition a QAM constellation into annular subconstellations of equal size by scaling a basic circular region, and we realize shaping gain through a shaping code which selects the subconstellations with different probabilities. It is possible to achieve any desired fractional bit rate through this method of shaping on rings.
Calderbank and Ozarow [calderbank2002nonequiprobable] proposed non-equiprobable signaling for the Gaussian channel, and the combination of trellis coding and shaping was essential to the V.34 voiceband modem standard [forney1989multidimensional, ungerboeck1982channel]. Wireless communication employs LDPC codes rather than trellis codes, and Section III-C describes how we integrate shaping into the calculation of log-likelihood ratios (LLRs) necessary for decoding.
@@ II-A Channel estimation
In practice, systems are required to estimate the channel at the receiver. To estimate the channel, a reference symbol or pilot is transmitted. In this paper, we consider the Kronecker pilot pattern (also called the Type-2 demodulation reference signal (DMRS)) [3gpp_ts_38_211, 3gpp_ts_38_212]. In this setting pilots are placed on specific subcarriers in an OFDM symbol. The estimates at the pilot locations are obtained using a least squares estimator [nee_prasad_ofdm_2000] and to get the estimate at the rest of the locations a two-dimensional interpolation is carried out. For the two-dimensional interpolation, the (pre-computed) time and frequency covariance matrices are used, which are computed as:
@@
respectively [^2] Here, we consider the diagonal values when we compute the covariance matrices. This is because under fractional delay-Doppler channel, the channel matrix is not strictly diagonal due to inter-carrier interference (ICI) in OFDM. Taking the whole matrix introduces ICI effects into the time and frequency covariance matrix.. A linear minimum mean squared error (LMMSE) interpolator is used for the two-dimensional interpolation of the channel values [hoeher1997two-dimensional] which gives the estimate of the diagonals of $\mathbf{H}_{\mathsf{F}}$. This forms the estimated channel matrix which is used for equalization.
@@ III-A Shaping Encoder
At the transmitter, an augmented constellation is generated. An example is shown for 4-QAM, in Fig. 2. The augmented constellation is a combination of an inner (or the original) constellation and an outer constellation [^3] In this paper, although we divide the constellation into two regions, inner and outer, the shaping idea is very general. For example, the constellation could be divided into multiple partitions as shown in [calderbank2002nonequiprobable].. Each symbol from the inner constellation is mapped to a point on the outer constellation. Such pairs have the same bit representation. In the example, if the point $1+1j$ is represented by the binary vector $0\ 0]^{\top}$ then so is the point 3+1j-3+1j. In the Figure, each constellation point on the 4-QAM shown by a square, circle, triangle, or star, is mapped to a point on the outer constellation represented by the corresponding shaded shape. The choice between the inner and outer constellation is provided by the shaping code. Note that the shaping code is not actually transmitted, the extra information bits are conveyed through power during transmission.
@@ III-B Shaping Decoder
However, since the shaping codeword has a pre-defined structure (for example, sparsity) the decoded codeword may not always be a valid codeword. In such cases, the index of the dirtiest symbol is computed as:
@@
and the bit at $\hat{\mathbf{c}}[i]$ is flipped. If this results in a valid codeword, the process is stopped else the process is continued to find the next dirtiest symbol index and so on. This procedure gives the decoded shaping codeword.
@@ III-C LDPC Decoding
LLR computation for LDPC decoding is presented in Algorithm 1. The algorithm takes as input the equalized symbols, the constellation points, the label map which corresponds to the labels for both inner and outer constellation (each label is mapped to exactly two symbols), noise variance, and the prior probabilities of the constellation points in the log scale. From Steps 3 – 9, a minimum distance metric is computed and the prior probabilities are introduced. From Steps 10 – 15, the Algorithm combines the metrics for all the points that have the same label using the log sum exponential approximation for improving numerical stability [bishop2006pattern]. Combining metrics is essential since more than one constellation point is mapped to the same index (see Sec. III-A). Steps 16 through 22 compute the logits to LLR conversion in the usual way.
@@ III-D Rate, PAPR, and Complexity Considerations
In practice, PAPR is an important consideration. A high PAPR requires an amplifier with large linear operating region [nee_prasad_ofdm_2000], which may not always be realizable in practice. When the constellation is augmented, we increase the peak power of the constellation. The average energy of the augemented constellation is:
@@ IV Numerical Results
In this Section we present the numerical results evaluating the performance of the OFDM scheme with shaping. For all the results presented here, we consider a practical setting where estimated channel is used (per Sec. II-A) for equalization along with LDPC encoding. We consider an OFDM system with $M=72,N=14,\Delta f=30$ kHz. For all the simulations we consider the Vehicular-A channel model [veh_a] with a maximum Doppler spread $\nu_{\max}$ of $815$ Hz. Each path Doppler is obtained using the Jake’s spectrum $\nu_{p}=\nu_{\max}\cos(\theta)$, where $\theta\sim\mathsf{U}[-\pi,\pi)$.
@@ IV-A $4$-QAM, $15$-length shaping
In this Subsection, we consider a $15$-length shaping code with sparsity $1$. We have $16$ codewords in the shaping codebook which convey $4$ bits of information over $15$ channel uses. The additional rate is $\approx 0.25$. We pair this with an LDPC code of rate $0.5$, and the effective rate is $1.25$. For the unshaped performance we consider a rate $5/8$ LDPC code with $4$-QAM that also achieves rate $1.25$. Figure 4 compares the bit-error performance of the two systems. It is seen that the performance of the OFDM system with shaping is superior compared to that without shaping and the performance gap increases with increase in signal power. Shaping gain [calderbank2002nonequiprobable] and lower rate for LDPC with shaping both contribute to this performance improvement.
@@ IV-B $4$-QAM, $23$-length shaping
In this Subsection, we consider a $23$-length shaping code with sparsity $3$. We have $2048$ codewords in the shaping codebook which convey $11$ bits of information over $23$ channel uses. The additional rate is $\approx 0.5$. We pair this with an LDPC code of rate $0.5$, and the effective rate is $1.5$. For the unshaped performance we consider a rate $3/4$ LDPC code with $4$-QAM that also achieves rate $1.5$. Figure 5 shows the bit-error performance of the shaped and unshaped OFDM systems. The performance gap is wider here, because to match the rate, the unshaped OFDM system uses a higher rate LDPC, which is less capable of correcting errors.
@@ IV-C $16$-QAM, $23$-length shaping
In this Subsection, we consider the same $23$-length shaping code. The effective rate is $2.5$ since we use $16$-QAM. For the unshaped performance we consider a rate $5/8$ LDPC code with $16$-QAM that also achieves rate $2.5$. Figure 6 shows the bit-error comparison. At high signal energies there is performance gain of about $1$ dB over the unshaped OFDM system.

### v2

# Increasing Data Rate through Shaping on Wireless Channels Subject to Mobility and Delay Spread

@@ Abstract
This letter describes how to improve performance of cellular systems by combining non-equiprobable signaling with low-density parity check (LDPC) coding for an orthogonal frequency division multiplexing system. We focus on improving performance of wireless transmission with rate $3/4$ LDPC codes employing $4$-QAM at the cell edge. We double the size of the $4$-QAM constellation by introducing a second shell of signal points, and we implement non-equiprobable signaling through a shaping code which selects the high energy shell less frequently than the low energy shell. We emphasize that the shaping code is also used to transmit information. We employ rate $1/2$ LDPC coding and select the rate of the shaping code to match that of rate $3/4$ LDPC coding using $4$-QAM. We describe how to combine coding and shaping by integrating shaping into the calculation of log-likelihood ratios (LLRs) necessary for decoding LDPC codes. We present simulation results for a representative Veh-A channel showing gains of $4$ dB at a bit error rate (BER) of $10^{-3}$. Our results suggest that it is better to increase rate at the cell edge by introducing shaping, rather than by increasing the rate of the LDPC code.
@@ I Introduction
In his 1948 paper, Shannon [10] recognized that signals input to a Gaussian channel should themselves be selected with a Gaussian distribution. Wireless channels subject to mobility and delay spread are not Gaussian channels, but it is still advantageous to select signals with large energy less frequently than signals with small energy. This can be achieved by geometric shaping (GS), where a standard QAM constellation is warped so that points close to the origin are closer together and points closer to the periphery are further apart. A second method is probabilistic shaping (PS), where the constellation points are kept as is, and the distribution over them is changed. These shaping methods apply to any coding and modulation scheme, they do not introduce additional rate, and they enable modest improvements in performance (typically less than 1dB) [8].
In their 1998 survey paper, Forney and Ungerboeck [5] described how coding and shaping can be combined to allow capacity to be approached on any linear Gaussian channel. A particular focus is the 1994/1996 ITU-T V.34 standard that enabled wireline access to the internet at rates up to $33.6$ kb/s. the survey describes how error performance on the Gaussian channel is determined by the figure of merit $d^{2}/P$ where $P$ is the average transmitted signal power and $d^{2}$ is the minimum squared distance of the code. We can transmit an extra bit by doubling the size of a baseline QAM constellation and signaling equiprobably. However, this doubles the average transmitted signal power $P$, and to maintain error performance we would need to double the minimum squared distance $d^{2}$. In the 1990s computing power was very limited and the complexity of decoding the more powerful code was typically out of reach. The V.34 standard increased transmission rate without increasing coding complexity by using a rate $R$ shaping code to transmit additional information. The V.34 standard includes the method of shaping on rings proposed by Calderbank and Ozarow [4] where a QAM constellation is partitioned into annular subconstellations of equal size by scaling a basic circular region, and the shaping code selects the subconstellations with different probabilities. Equiprobable signaling would match the rate by scaling the baseline constellation by $2^{R}$, thereby increasing average transmitted signal power by the same factor. When the power required by non-equiprobable signaling is less than $2^{R}P$ there is a shaping gain. It is possible to closely approach the asymptotic shaping gain of $1.53$ dB on a Gaussian channel by shaping on $8$ annuli (see [4] for details).
We use the method of shaping on rings to improve 5GNR performance at the cell edge. The 5GNR standard specifies rate $1/2$ and rate $3/4$ LDPC codes which use 4-QAM at the cell edge. At a bit error rate of $10^{-3}$ the gap between the performance of the rate $1/2$ code and the rate $3/4$ code is significant - approximately $4$ dB. We bridge this gap by combining the rate $1/2$ LDPC code with a shaping code. We expand the baseline $4$-QAM constellation to the $8$-QAM constellation [^1] We refer to the expanded constellation as $8$-QAM since the expanded constellation has $8$ points. The constellation can also be viewed as a $16$-QAM constellation with the points not present in Fig. 2 carrying zero probability. shown in Fig. 2 by introducing a second shell of signal points, and we employ shaping codes that select the high energy shell less frequently than the low energy shell. The rate $3/4$ LDPC code transmits $1.5$ bits per QAM symbol, and we match this rate by introducing a shaping code with rate $0.5$ bits per QAM symbol. We ask the fundamental question of whether it is better to increase rate at the cell edge in 5GNR by introducing shaping or by increasing the rate of the LDPC code. We present simulation results for a representative Veh-A channel showing that rate $1/2$ LDPC coding with shaping improves on rate $3/4$ LDPC coding by $4$ dB at a bit error rate (BER) of $10^{-3}$. Our results demonstrate that at the cell edge where transmission rates are low, there is significant value in using shaping to contribute transmission rate. This does not appear possible with the GS and PS methods currently under consideration at 3GPP.
In order to combine coding and shaping it is necessary to integrate shaping into the calculation of the log-likelihood ratios (LLRs) necessary for decoding LDPC codes. We consider the orthogonal frequency division multiplexing (OFDM) system model because that is the current focus of 3GPP, but the method of shaping on rings can be applied to any modulation, for example to orthogonal time frequency space (OTFS). Section III presents our algorithm for calculating LLRs. A feature of this algorithm that may be of independent interest is the way it takes signal energy into account. The gains from our method of shaping diminish as the transmission rate increases, but we are still able to present an example in Section IV-C showing significant gains for $16$-QAM.
@@ II-A Channel estimation
In practice, systems are required to estimate the channel at the receiver. To estimate the channel, a reference symbol or pilot is transmitted. In this paper, we consider the Kronecker pilot pattern (also called the Type-2 demodulation reference signal (DMRS)) [1, 2]. In this setting pilots are placed on specific subcarriers in an OFDM symbol. The estimates at the pilot locations are obtained using a least squares estimator [11] and to get the estimate at the rest of the locations a two-dimensional interpolation is carried out. For the two-dimensional interpolation, the (pre-computed) time and frequency covariance matrices are used, which are computed as:
@@
respectively [^3] Here, we consider the diagonal values when we compute the covariance matrices. This is because under fractional delay-Doppler channel, the channel matrix is not strictly diagonal due to inter-carrier interference (ICI) in OFDM. Taking the whole matrix introduces ICI effects into the time and frequency covariance matrix.. A linear minimum mean squared error (LMMSE) interpolator is used for the two-dimensional interpolation of the channel values [6] which gives the estimate of the diagonals of $\mathbf{H}_{\mathsf{F}}$. This forms the estimated channel matrix which is used for equalization.
@@ III-A Shaping Encoder
At the transmitter, an augmented constellation is generated. An example is shown for 4-QAM, in Fig. 2. The augmented constellation is a combination of an inner (or the original) constellation and an outer constellation [^4] In this paper, although we divide the constellation into two regions, inner and outer, the shaping idea is very general. For example, the constellation could be divided into multiple partitions as shown in [4].. Each symbol from the inner constellation is mapped to a point on the outer constellation. Such pairs have the same bit representation. In the example, if the point $1+1j$ is represented by the binary vector $0\ 0]^{\top}$ then so is the point 3+1j-3+1j. In the Figure, each constellation point on the 4-QAM shown by a square, circle, triangle, or star, is mapped to a point on the outer constellation represented by the corresponding shaded shape. The choice between the inner and outer constellation is provided by the shaping code. Note that the shaping code is not actually transmitted, the extra information bits are conveyed through power during transmission.
@@ III-B Shaping Decoder
However, since the shaping codeword has a pre-defined structure (for example, sparsity) the decoded codeword may not always be a valid codeword. To get to the valid codeword, we use the fact that while decoding constellation symbols, received symbols closer to the decision boundary are less reliable than the ones closer to the constellation points. The less reliable received symbols need to be moved a greater distance to be decoded to the nearest codeword compared to the more reliable ones. The less-reliable symbols are therefore more likely to be in error. This motivates the following decoding rule. When the codeword is not valid, the index of the “dirtiest” symbol (i.e., the symbol that needs to be moved a great distance for decoding) is computed as:
@@
and the bit at $\hat{\mathbf{c}}[i]$ is flipped (since it is the least reliable). If this results in a valid codeword, the process is stopped else the process is continued to find the next dirtiest symbol index and so on. This procedure gives the decoded shaping codeword.
@@ III-C LDPC Decoding
LLR computation for LDPC decoding is presented in Algorithm 1. The algorithm takes as input the equalized symbols, the constellation points, the label map which corresponds to the labels for both inner and outer constellation (each label is mapped to exactly two symbols), noise variance, and the prior probabilities of the constellation points in the log scale. From Steps 3 – 9, a minimum distance metric is computed and the prior probabilities are introduced [^5] The probability of $1$s in the shaping code provides the priors for the outer constellation points and the probability of 0s provides the priors for inner constellation points. This is because once the choice is made between the inner and outer constellation, the information symbols are chosen uniformly at random.. From Steps 10 – 15, the Algorithm combines the metrics for all the points that have the same label using the log sum exponential approximation for improving numerical stability [3]. Combining metrics is essential since more than one constellation point is mapped to the same index (see Sec. III-A). Steps 16 through 22 compute the logits to LLR conversion in the usual way.
@@ III-D Rate, PAPR, and Complexity Considerations
In practice, PAPR is an important consideration. A high PAPR requires an amplifier with large linear operating region [11], which may not always be realizable in practice. When the constellation is augmented, we increase the peak power of the constellation. The average energy of the augemented constellation is:
@@ IV Numerical Results
In this Section we present the numerical results evaluating the performance of the OFDM scheme with shaping. For all the results presented here, we consider a practical setting where estimated channel is used (per Sec. II-A) for equalization along with LDPC encoding. We consider an OFDM system with $M=72,N=14,\Delta f=30$ kHz. For all the simulations we consider the Vehicular-A channel model [9] with a maximum Doppler spread $\nu_{\max}$ of $815$ Hz. Each path Doppler is obtained using the Jake’s spectrum $\nu_{p}=\nu_{\max}\cos(\theta)$, where $\theta\sim\mathsf{U}[-\pi,\pi)$. We use the 5G LDPC encoder and decoder provided by Sionna [7]. The message length $k$ and code length $n$ are specified for each code in the following subsections.
@@ IV-A $4$-QAM, $15$-length shaping
In this Subsection, we consider a $15$-length shaping code with sparsity $1$. We have $16$ codewords in the shaping codebook which convey $4$ bits of information over $15$ channel uses. The additional rate is $\approx 0.25$. We pair this with an LDPC code of rate $0.5$ ($k=855,n=1710$), and the effective rate is $1.25$. For the unshaped performance we consider a rate $5/8$ ($k=855,n=1710$) LDPC code with $4$-QAM that also achieves rate $1.25$. Figure 4 compares the bit-error performance of the two systems. It is seen that the performance of the OFDM system with shaping is superior compared to that without shaping and the performance gap increases with increase in signal power. Shaping gain [4] and lower rate for LDPC with shaping both contribute to this performance improvement.
@@ IV-B $4$-QAM, $23$-length shaping
In this Subsection, we consider a $23$-length shaping code with sparsity $3$. We have $2048$ codewords in the shaping codebook which convey $11$ bits of information over $23$ channel uses. The additional rate is $\approx 0.5$. We pair this with an LDPC code of rate $0.5$ ($k=851,n=1702$), and the effective rate is $1.5$. For the unshaped performance we consider a rate $3/4$ ($k=1296,n=1728$) LDPC code with $4$-QAM that also achieves rate $1.5$. Figure 5 shows the bit-error performance of the shaped and unshaped OFDM systems. The performance gap is wider here, because to match the rate, the unshaped OFDM system uses a higher rate LDPC, which is less capable of correcting errors.
@@ IV-C $16$-QAM, $23$-length shaping
In this Subsection, we consider the same $23$-length shaping code with LDPC code of rate $0.5$ ($k=1702,n=3404$). The effective rate is $2.5$ since we use $16$-QAM. For the unshaped performance we consider a rate $5/8$ ($k=2160,n=3456$) LDPC code with $16$-QAM that also achieves rate $2.5$. Figure 6 shows the bit-error comparison. At high signal energies there is performance gain of about $1$ dB over the unshaped OFDM system.
@@
Figure 7: CCDF of PAPR comparing the performance of OFDM with no shaping, $15$-length shaping and $23$-length shaping for 4-QAM constellation.
IV-D PAPR
Figure 7 shows the PAPR of the OFDM system with and without shaping. 4-QAM constellation is used for communication. It is seen that although the PAPR of the OFDM system is expected to increase when the information symbols are picked from the outer constellation, after the IDFT block at the transmitter, the difference in PAPR between the systems with and without shaping is negligible. This is because of the sparsity of the shaping codeword and also the fact that DFT bases spread the energy over all time domain symbols.